OpenStudy (anonymous):

The equatorial radius of Earth is approximately 6 × 103 km, while the equatorial radius of Saturn is approximately 6 × 104 km. Which of the following is true? A. The equatorial radius of Saturn is approximately one hundred times that of Earth. B. The equatorial radius of Earth is approximately one hundred times that of Saturn. C. The equatorial radius of Earth is approximately ten times that of Saturn. D. The equatorial radius of Saturn is approximately ten times that of Earth.

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@LannaBoo24

OpenStudy (anonymous):

@SiennaSurfer @professional_loser_

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

i think a

OpenStudy (anonymous):

ok thanksssssss soooooo much

OpenStudy (anonymous):

can u write something so i can give u a medal

OpenStudy (anonymous):

\[\frac{ 10^4 }{10^3 }=10^4*10^{-3}=10^{4-3}=10\] what can you say now?

OpenStudy (anonymous):

you got it.

OpenStudy (anonymous):

i am not helping you for medal. can you say what is your option ?